I have bean asked to consider in detail the question x2 =-5x-6
I know how to solve the equation, however I have a feeling I could be asked why you put it equal to zero? Could you please tell me the reason along with any answers to any other questions you think I may be asked.?It's for an interview for teacher training. Thankyou

Answer by abhinav75(8)

it is not essential to put it as zero, but it is more convenient..
for example, x^2 = -5x -6
=> x^2 + 5x + 6=0
=> (x+2)(x+3) = 0
x= -2,-3
now, let us take this case..
x^2 = - 5x -6
=> x^2 + 5x + 4 = -2
=> (x+4)(x+1) = -2
now, -2 can be factorised as (-1*2) or (-2*1)
so, case 1..
-2 = -1 *2
so, x+1 = -1, x+4 = 2
both give x = -2
similarly, in second case, we get x = -3...
its all a matter of convenience..